OpenStudy (anonymous):

Lamar invested his savings in two investment funds. The amount he invested in Fund A was 4 times as much as the amount he invested in Fund B. Fund A returned a 3% profit and Fund B returned a 6% profit. How much did he invest in Fund B, if the total profit from the two funds together was $1620?

OpenStudy (anonymous):

If we use \[ A = \ \textrm{amount invested in Fund A}\] \[ B = \ \textrm{amount invested in Fund B}\] \[P_A = \textrm{Profit on Fund A}\] \[P_B = \textrm{Profit on Fund B}\] How are each of those variables related?

OpenStudy (anonymous):

In the problem it says that the amount invested in Fund A was 4 times the amount invested in Fun B, so we can say \[ A = 4B\] The problem also says that the Profit on Fund A was 3%. 3% = 3/100 so we can say \[P_A = \frac{3}{100}A\] Profit on Fun B was 6%. \[P_B = \frac{6}{100}B\] And lastly, that total Profits were $1620 \[P_A + P_B = $1620\] How are P_A and P_B related to A and B? And how is A related to B?